Skip to content
DeepForgeby svx
All problems
MediumProbabilitypr-035

Card Draw At Least One

Problem

Probability of drawing at least one target card in draws cards taken without replacement from a deck.

P(at least one) = 1 - C(deck - targets, draws) / C(deck, draws). Return 0.0 when deck, targets or draws is not positive and 1.0 when every draw must be a target.

Hint

Complement of drawing only non-target cards.

Test cases

case 1
in: [52,4,5]
expected: 0.34115800166220334
case 2
in: [52,4,0]
expected: 0
case 3
in: [10,10,3]
expected: 1
case 4
in: [6,1,6]
expected: 1
case 5
in: [10,0,5]
expected: 0
Starter codePython
import math


def card_draw_at_least_one(deck, targets, draws):
    # Your code here
    pass

5 test cases · run your code against them in the editor